OpenStudy (anonymous):

The sum and product of the roots of a quadratic equation are -2/5 and -1/10, respectively. What is the equation? A.5x^2+2x-2=0 B.5x^2-2x+2=0 C.10x^2+4x-1=0 D.10x^2-4x+1=0

OpenStudy (perl):

vieta's formula says for the equation ax^2 + bx + c =0 with roots r1,r2 r1 + r2 = -b/a r1*r2 = c/a

OpenStudy (anonymous):

thank you